OpenStudy (anonymous):

prove csc^4x-cot^4x=2csc^2x-1

OpenStudy (anonymous):

csc^4x-cot^4x = 2csc^2x-1 csc^4(x) - (cot^2(x))^2 = 2csc^2x-1 csc^4(x) - (csc^2(x) - 1)^2 = 2csc^2x-1 csc^4(x) - (csc^4(x) - 2csc^2(x) +1) = 2csc^2x-1 csc^4(x) - csc^4(x) + 2csc^2(x) -1 = 2csc^2x-1 2csc^2(x) - 1 = 2csc^2x-1 Medal O_O

OpenStudy (anonymous):

Thank you!